OpenStudy (anonymous):

Choose the correct simplification of (7x^3 - 8x - 5) + (3x^3 + 7x + 1). Answer 10x^3 + x + 4 4x^3 + 15x + 6 4x^3 - 15x - 6 10x^3 - x - 4

OpenStudy (hoblos):

combine like terms... =7x^3 - 8x - 5 +3x^3 + 7x + 1 = (7+3)x^3 +(-8+7)x + (-5+1) = 10x^3 - x - 4